# MEV Capture Through Time-Advantaged Arbitrage

### v1

@@ 3.3 Simulation Results
We run the simulations using the following parameters, which are intended for implementation on Arbitrum: $\Delta=10$ ms, $T_{w}=200$ ms, $T=1$ minute. [^4] Experimenting with various parameter values did not produce significantly different results.
@@
We simulate the optimal strategy for this type of price distribution, assuming zero drift and a daily volatility ranging from 1% to 10%. Our results show that the best strategy for the time-advantaged arbitrageur is to wait for the entire duration of their time advantage. Specifically, for any price difference $p$ where $profit(p)>0$ and for $t_{w}<T_{W}$ and $t<T$, we find that $wait(p,t_{w},t)>arb(p,t_{w},t)$, i.e., the expected profit from waiting exceeds that from executing the arbitrage immediately.
@@
We gathered one month of bid and ask prices for various trading pairs from March 2024.555We consider perpetual futures prices since Binance – the largest cryptocurrency exchange – offers full historical data for their top order book levels while only providing second-by-second price data for spot pairs. Perpetual futures often have higher trading volumes compared to spot markets, emphasizing that their prices are likely to accurately reflect underlying asset prices. Using this data, we compute the historical distribution of the change in the mid-price over a time of $\Delta=10$ ms.
@@
We perform this analysis for several trading pairs, including Ethereum (ETH-USDT) and Shiba Inu (SHIB-USDT). In each case, we consistently find that waiting for the duration of the arbitrageur’s time advantage yields the optimal strategy.
@@
4 Arbitrage Profits
First-Come First-Serve (FCFS).
Priority Gas Auction (PGA).
FCFSwith a time-advantaged actor (Timeboost).
4.1 Simulated Results
@@
Note that the analysis deriving the optimal strategy under Timeboost assumes independently distributed price changes. However, when prices exhibit mean reversion — i.e., negative autocorrelation — profits under FCFScan match or even exceed those from a waiting strategy with Timeboost, as seen with the ETH-BTC and ARB-ETH pairs. In cases of negative autocorrelation, where it is more likely that a price move that opens an arbitrage opportunity is followed by a reversal, it becomes more advantageous to exploit arbitrage opportunities immediately.
@@ 4.2 Capturing Arbitrage Profits for AMMs under the Timeboost Policy
Attributing MEVand returning (part of) it to the users is an actively studied area, with multiple theoretical and practical proposals, such as Adaptive Curves [11] and MEVTaxes [13]. Under a time-advantage policy (such as Timeboost), one potential general approach to MEV capture would involve separately auctioning time advantages for interacting with specific contract addresses, such as AMM pools. The proceeds from each auction could then be returned to the relevant contract.
@@
In the following, we analyze how much arbitrage profit AMMs could capture through this approach. Without any adaptation, the arbitrageur’s profit from moving the pool’s marginal price from $P_{0}$ to the true price $P$, given pool reserves of $(X,Y)$, equals $profit(P_{0},P):=XP-2\sqrt{XYP}+Y$. We assume the pool attempts to capture a $\alpha$ fraction of this profit. In other words, the AMM will offer a trade that shifts its marginal price to $P$ while incurring a loss of $(1-\alpha)profit(P_{0},P)$, for any $P$. Hence, it allows for its reserves to be changed to $(X_{1},Y_{1})$, such that $\frac{Y_{1}}{X_{1}}=P$, and $(X-X_{1})P-(Y_{1}-Y)=(1-\alpha)profit(P_{0},P)$. This fully determines the AMM’s trading behavior for time-advantaged transactions. The AMM allows changing its reserves to:
@@ Appendix 0.A Arbitrage Profits for Constant Product Pools
Consider a CPMMpool with trading fee $f$. In the following, we calculate the profit of the maximal arbitrage trade, that is, the trade that produces the maximum arbitrage profit, depending on the price difference between the AMM pool and the external market. Recall that we assume the price impact of the arbitrage trade on the external market to be negligible.
@@
Next, we consider the relative change in the pool’s value resulting from the maximal arbitrage trade executed at the price difference $p$. For $P>\frac{1}{1-f}P_{0}$, the pool reserves change $(X,Y)$ to $\left(\sqrt{\frac{XY}{(1-f)P}},\sqrt{XY(1-f)P}\right)$. In particular, the pool’s value at the respective pool price changes from $2Y$ to $2\sqrt{XY(1-f)P}$, i.e. by a factor of $\sqrt{1-f}\frac{P}{P_{0}}$. Applying the same reasoning to the case where $P<(1-f)P_{0}$ results in

### v2

@@ 3.3 Simulation Results
We run the simulations using the following parameters, which are intended for implementation on Arbitrum: [^3] Experimenting with various parameter values did not produce significantly different results. $\Delta=10$ ms, $T_{w}=200$ ms, $T=1$ min, and consider commonly used values for the pool’s trading fee, including 0.05% and 0.3%. For the price of the volatile token, we analyze the following two scenarios.
@@
We simulate the optimal strategy for this type of price distribution, assuming zero drift and a daily volatility ranging from 1% to 10%. Our results show that the best strategy for the time-advantaged arbitrageur is to wait for the entire duration of their time advantage. Specifically, for any relative price difference $p$ where $profit(p)>0$ and for $t_{w}<T_{w}$ and $t<T$, we find that $wait(p,t_{w},t)>arb(p,t_{w},t)$, i.e., the expected profit from waiting exceeds that from executing the arbitrage immediately.
@@
We gathered one month of bid and ask prices for various trading pairs from March 2024.444We consider perpetual futures prices since Binance – the largest cryptocurrency exchange – offers full historical data for their top order book levels while only providing second-by-second price data for spot pairs. Perpetual futures often have higher trading volumes compared to spot markets, emphasizing that their prices are likely to accurately reflect underlying asset prices. Using this data, we compute the historical distribution of the change in the mid-price over a time of $\Delta=10$ ms.
@@
We perform this analysis for several trading pairs, including Ethereum (ETH-USDT) and Shiba Inu (SHIB-USDT). In each case, our analysis consistently demonstrates that the optimal strategy is to wait for the full duration of the arbitrageur’s time advantage.
4 Arbitrage Profits
First-Come First-Serve (FCFS).
Priority Gas Auction (PGA).
FCFSwith a time-advantaged actor (Timeboost).
@@ 4 Arbitrage Profits
4.1 Simulation Results
@@
Note that the analysis deriving the optimal strategy under Timeboost assumes independently distributed price changes. However, when prices exhibit mean reversion – i.e., negative autocorrelation – profits under FCFScan match or even exceed those from a waiting strategy with Timeboost, as seen with the ETH-BTC and ARB-ETH pairs. In cases of negative autocorrelation, where it is more likely that a price move that opens an arbitrage opportunity is followed by a reversal, it becomes more advantageous to exploit arbitrage opportunities immediately.
@@ 4.2 Capturing Arbitrage Profits for AMMs under the Timeboost Policy
Attributing MEVand returning (part of) it to users is an actively studied area, with multiple theoretical and practical proposals, such as Adaptive Curves [12] and MEVTaxes [14]. Under a time-advantage policy (such as Timeboost), one potential general approach to MEV capture would involve separately auctioning time advantages for interacting with specific contract addresses, such as AMM pools. The proceeds from each auction could then be returned to the relevant contract.
@@
In the following, we analyze how much arbitrage profit AMMs could capture through this approach. Without any adaptation, the arbitrageur’s profit from moving the pool’s marginal price from $P_{0}$ to the true price $P$, given pool reserves of $(X,Y)$, equals $profit(P_{0},P):=XP-2\sqrt{XYP}+Y$. We assume the pool attempts to capture a $\alpha$ fraction of this profit. In other words, the AMM will offer a trade that shifts its marginal price to $P$ while incurring a loss of $(1-\alpha)profit(P_{0},P)$, for any $P$. Hence, it allows its reserves to be changed to $(X_{1},Y_{1})$, such that $\frac{Y_{1}}{X_{1}}=P$, and $(X-X_{1})P-(Y_{1}-Y)=(1-\alpha)profit(P_{0},P)$. This fully determines the AMM’s trading behavior for time-advantaged transactions. The AMM allows changing its reserves to:
@@ Appendix 0.A Arbitrage Profits for Constant Product Pools
Consider a CPMMpool with trading fee $f$. In the following, we calculate the profit of the maximal arbitrage trade, that is, the trade that produces the maximum arbitrage profit, depending on the relative price difference between the AMM pool and the external market. Recall that we assume the price impact of the arbitrage trade on the external market to be negligible.
@@
Next, we consider the relative change in the pool’s value resulting from the maximal arbitrage trade executed at the relative price difference $p$. For $P>\frac{1}{1-f}P_{0}$, the pool reserves change $(X,Y)$ to $\left(\sqrt{\frac{XY}{(1-f)P}},\sqrt{XY(1-f)P}\right)$. In particular, the pool’s value at the respective pool price changes from $2Y$ to $2\sqrt{XY(1-f)P}$, i.e. by a factor of $\sqrt{1-f}\sqrt{\frac{P}{P_{0}}}$. Applying the same reasoning to the case where $P<(1-f)P_{0}$ results in